OpenStudy (anonymous):

HELP PLEASE I WILL GIVE MEDAL! At a football game, a vender sold a combined total of 162 sodas and hot dogs. The number of sodas sold was 48 more than the number of hot dogs sold. Find the number of sodas sold and the number of hot dogs sold.

OpenStudy (anonymous):

s + h = 162 s = h + 48 now sub in h + 48 in for s in the first equation and solve for h. Once h is found, sub that back into either of the equations and find s.

OpenStudy (anonymous):

so its going to be 2h+48=162?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay so I got 57 after solving everything

OpenStudy (anonymous):

ok....h = 57. Now sub that back into either equation to find s

OpenStudy (anonymous):

so 57 hotdogs sold and 105 sodas sold?

OpenStudy (anonymous):

lets check it..... s + h = 162 105 + 57= 162 162 = 162 (correct) s = h + 48 105 = 57 + 48 105 = 105 (correct) yes....57 hotdogs and 105 sodas

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

no problem :)